# Scale-Dependent Multifractality in Bitcoin Realised Volatility: Implications for Rough Volatility Modelling

### v2

@@
Scale-Dependent Multifractality in Bitcoin Realised Volatility: Implications for Rough Volatility Modelling
@@
September 20, 2025
We assess the applicability of rough volatility models to Bitcoin realised volatility using the normalised $p$-variation framework of Cont and Das, (2024). Applying this model-free estimator to high-frequency Bitcoin data from 2017 to 2024 across multiple sampling resolutions, we find that the normalised statistic remains strictly negative throughout, precluding the estimation of a valid roughness index. Stationarity tests and robustness checks reveal no significant evidence of non-stationarity or structural breaks as explanatory factors. Instead, convergent evidence from three complementary diagnostics, namely Multifractal Detrended Fluctuation Analysis, log-log moment scaling, and wavelet leaders, reveals a multifractal structure in Bitcoin volatility. This scale-dependent behaviour violates the homogeneity assumptions underlying rough volatility estimation and accounts for the estimator’s systematic failure. These findings suggest that while rough volatility models perform well in traditional markets, they are structurally misaligned with the empirical features of Bitcoin volatility.
@@ 1.1 Background and Motivation
The rough volatility framework has emerged as a prominent approach for modelling the fine-scale dynamics of financial market volatility. This development is grounded in empirical findings that indicate that realised volatility paths possess low Hölder regularity, which implies irregularity in their sample paths at high frequencies. Rough stochastic volatility models have been shown to replicate empirical features commonly observed in financial markets, including volatility clustering, short-term persistence, and the skew and term structure of implied volatility surfaces in equity and index markets Gatheral et al., (2018); Fukasawa et al., (2022). These observations have motivated attempts to extend the applicability of rough volatility models to nontraditional asset classes.
However, the extent to which this framework generalises to structurally distinct markets remains an open empirical question. Bitcoin, the most liquid and actively traded cryptocurrency, offers a particularly informative out-of-sample test. Unlike traditional equity or FX markets, Bitcoin exhibits fragmented liquidity across exchanges Makarov and Schoar, (2020), frequent discontinuities due to jumps Lahmiri and Bekiros, (2020), regime-switching volatility Caporale and Zekokh, (2021), and an evolving market microstructure shaped by protocol changes and shifting incentives Easley et al., (2019). This motivates a closer empirical examination of whether rough volatility assumptions hold in such a market.
@@
Prior applications of rough volatility methodologies to Bitcoin have predominantly employed parametric inference techniques or heuristic scaling diagnostics, often without systematically assessing assumptions of homogeneous Hölder continuity Bianchi et al., (2022). Empirical studies on cryptocurrencies more broadly, including Liu and Tsyvinski, (2021), have concentrated on asset pricing and factor models, without directly addressing the scaling properties of volatility. The normalised $p$-variation estimator recently introduced by Cont and Das, (2024) provides a model-free approach to estimating the pathwise roughness of volatility signals. This estimator offers two notable advantages over traditional methods based on log-volatility increments or variogram regressions: (i) it exhibits improved asymptotic properties under general conditions, and (ii) it is less sensitive to microstructure noise and finite-sample distortions. These properties make it especially suitable for high-frequency volatility data in contexts like cryptocurrency markets, where noise and irregular sampling are common.
This study builds on the normalised $p$-variation estimator introduced by Cont and Das, (2024) to perform a systematic evaluation of the roughness of Bitcoin realised volatility across a range of sampling frequencies and time horizons. This analysis examines the applicability of rough volatility models to Bitcoin and contributes to a deeper understanding of the structural features that differentiate cryptocurrency volatility dynamics.
@@ 1.2 Research Gap
Bitcoin’s volatility displays structural features such as multiscale clustering, episodic jumps, and irregular liquidity that differ markedly from the conditions under which rough volatility models have been validated.For instance, Bianchi et al. (2022) report low Hurst exponents for Bitcoin by fitting rough stochastic volatility models and performing scaling regressions on realised volatility. However, their approach implicitly assumes a global scaling regime and stationarity, without subjecting these conditions to empirical verification. These assumptions are critical, and their violation under multifractal dynamics or structural irregularities undermines the reliability of such roughness estimates. Crucially, no prior work has systematically examined whether Bitcoin realised volatility meets the homogeneity and stationarity conditions required for consistent estimation via the normalised $p$-variation method.
Moreover, recent research indicates that the observed roughness of realised volatility may arise from estimation error rather than genuine features of the spot volatility process Cont and Das, (2024); Rogers, (2019). In the case of Bitcoin, where multifractality and non-stationarity are theoretically plausible features, this raises serious doubts about the robustness and interpretive value of roughness estimates.
To our knowledge, no prior empirical study has systematically applied the $p$-variation diagnostic to Bitcoin realised volatility, nor has tested its structural compatibility with the assumptions underlying rough volatility models. It remains empirically unestablished whether Bitcoin volatility admits a well-defined roughness exponent. This study addresses this gap by applying the model-free normalised $p$-variation estimator alongside complementary diagnostic methods.
@@ 1.3 Research Question and Empirical Puzzle
This study tests whether Bitcoin realised volatility satisfies the homogeneous scaling assumptions necessary for consistent roughness estimation under the rough volatility paradigm.
We investigate this question by applying the normalised $p$-variation methodology of Cont and Das, (2024) to Bitcoin realised volatility across multiple years and sampling frequencies. Preliminary results indicate that, in contrast to equity and foreign exchange markets, Bitcoin volatility does not appear to admit a well-defined roughness index. This stands in contrast to prior findings for equity and foreign exchange markets, where roughness estimates have exhibited consistent behaviour and robust interpretability.
@@ 1.5 Contributions
First, it offers an extensive application of the normalised $p$-variation framework to Bitcoin realised volatility across multiple years and sampling frequencies, making the analysis more comprehensive than any prior roughness estimation in cryptocurrency markets. The analysis reveals a systematic and robust failure of roughness estimation across all empirical configurations.
@@
Third, the results contribute to the evaluation of the rough volatility paradigm by showing that, while effective in equity and foreign exchange markets, rough stochastic volatility models face structural challenges when applied to Bitcoin realised volatility. This finding underscores the need for models that explicitly incorporate multifractal and scale-dependent dynamics to capture the unique characteristics of cryptocurrency markets. Examples of such alternatives include the Multifractal Random Walk Bacry et al., (2001) and Multifractal Stochastic Volatility Calvet and Fisher, (2002) models.
@@ 2.1 Rough Volatility and the Hurst Paradigm
Empirical support for the rough volatility hypothesis has emerged from multiple methodological approaches. These include the log-regression of realised volatility increments Gatheral et al., (2018), parametric estimation of fractional stochastic volatility models Fukasawa et al., (2022), and non-parametric diagnostics based on $p$-variation Cont and Das, (2024). Across these methods, estimated Hurst exponents in equity and index markets typically fall in the range 0.1 to 0.3. These findings have motivated the adoption of rough stochastic volatility models, originally proposed by Gatheral and Rosenbaum and empirically validated by Bennedsen, Lunde, and Pakkanen Bennedsen et al., (2016), in both academic research and quantitative finance practice. Rough models have demonstrated empirical success in capturing the fine-scale behaviour of implied volatility surfaces and short-term volatility dynamics.
The applicability of the rough volatility paradigm across asset classes remains contested. Rogers, (2019) and Cont and Das, (2024) have emphasised that low roughness estimates derived from realised volatility may not reflect intrinsic properties of the spot volatility process, but instead result from structural features of the data or estimation artefacts. For example, Rogers, (2019) showed that microstructure noise or irregular sampling can generate artificially low Hurst exponents, cautioning against interpreting realised path roughness as direct evidence of true volatility roughness. These confounding effects, including multifractal scaling, may induce apparent roughness even when the underlying volatility process is not genuinely rough. This underscores the need for model-free diagnostics when extending rough volatility models beyond traditional markets.
@@ 2.2 The Cont–Das Normalised $p$-Variation Estimator
To address the limitations of traditional approaches, we adopt the framework introduced by Cont and Das, (2024). This framework provides a model-free, pathwise-consistent estimator of roughness based on the scaling behaviour of the normalised $p$-variation statistic $W(L,K,p)$, computed over a sequence of block partitions of the observed time series. Intuitively, $W(L,K,p)$ measures how the $p$-th variation of the process accumulates across blocks, normalised in such a way that for a self-similar process with a global scaling law, the statistic equals one exactly at the correct scaling exponent $p^{*}$. Thus, plotting $\log W(L,K,p)$ as a function of $1/p$, one expects a zero-crossing at a unique value $p=p^{*}$, which defines the variation index and yields the corresponding roughness index $H=1/p^{*}$. The existence of this zero-crossing serves as a rigorous criterion for determining whether the process admits a global monofractal scaling regime consistent with Hölder-type roughness. For reference, the formal definition of $W(L,K,p)$ is provided in Section 3.1.
The normalised $p$-variation estimator is non-parametric, robust to non-Gaussian features, and directly applicable to realised volatility series. It serves not only as a roughness estimator but also as a structural diagnostic of the underlying process. Specifically, the presence or absence of a zero-crossing in $\log W(L,K,p)$ determines whether the data admit a well-defined global scaling law consistent with the rough volatility paradigm. If no zero-crossing is observed, it implies that no single roughness exponent can characterise the process, indicating a fundamental incompatibility with monofractal scaling assumptions.
Recent applications of the method to equity markets have validated its efficacy Cont and Das, (2024), producing roughness estimates consistent with the existing literature. In particular, Cont and Das apply the normalised $p$-variation estimator to 5-minute realised volatility of the S&P 500 using the Oxford-Man Institute dataset. They observe a well-defined zero-crossing in $\log W(L,K,p)$, with the resulting roughness index estimated in the range $H\in[0.05,0.25]$, consistent with prior findings from log-regression methods.
@@ 2.3 Multifractality in Financial Time Series
Empirical research has established that financial time series often exhibit multifractal properties. Foundational models such as the Multifractal Random Walk Bacry et al., (2001) and the Multifractal Stochastic Volatility model Calvet and Fisher, (2002) account for stylised features such as volatility clustering, heavy tails, and long memory through underlying multifractal dynamics. Subsequent studies have confirmed these properties across equities, foreign exchange, and commodities. More recent work has extended this evidence to cryptocurrencies. High-frequency Bitcoin returns and trading volume exhibit multifractal scaling when analysed using Multifractal Detrended Fluctuation Analysis Mensi et al., (2019); Matteo, (2007), and complex multiscale dependence has been observed in cross-correlation structures Watorek et al., (2022). Additional research shows that Bitcoin realised volatility also displays a multifractal structure, though weaker than that of returns Zhou et al., (2022). These findings support the view that multifractality is a robust and persistent feature of financial markets and offer a plausible explanation for the structural deviations from monofractal scaling observed in Bitcoin volatility.
@@ 3.2 Data Description and Preprocessing
A 90% completeness threshold was imposed at each frequency to balance data quality with temporal coverage. For 1-minute and 5-minute data, all years from 2017 to 2024 met this criterion. For 10-minute data, 2022 and 2023 were excluded, while for 15-minute data, 2017, 2022, and 2023 were excluded. The remaining data were sufficient to observe clear patterns, and no contradictory behaviour was seen in partially excluded years at other frequencies. The complete list of retained and excluded years is given in Appendix Table A.1.
@@
Temporal structure and integrity of each retained window were further assessed. Large temporal gaps, defined as sequences of missing data exceeding one hour, were systematically identified at each frequency. Such gaps were rare for 1-minute and 5-minute data, but more common at coarser resolutions, especially in early years and in 2024. Full gap statistics are provided in Appendix Table A.2.
@@
3.3 Realised Volatility Construction
@@
This absolute return series preserves the temporal resolution of the data and captures the local magnitude of return fluctuations. For notational convenience, we refer to $\{RV_{t}\}$ as the realised volatility series, although it is more accurately a proxy for instantaneous volatility rather than an integrated measure.
@@ 3.4 Diagnostic Exploration: Roughness vs $K$
Prior to finalising the roughness estimation procedure, we conduct a diagnostic analysis to examine how the estimated roughness $H_{L,K}$ depends on the number of blocks $K$ (which controls the partition size). This step ensures we choose an appropriate $K$ for each case, as the estimate can be sensitive to this choice. The diagnostic analysis evaluates the behaviour of the roughness estimator $\widehat{H}^{L,K}$ as a function of the number of blocks $K$.
@@
To mitigate these issues, Cont and Das, (2024) recommend selecting $K$ such that $K\approx\sqrt{L}$, where $L$ is the number of available observations. This choice ensures a trade-off between bias and variance, grounded in the theoretical convergence properties of the normalised $p$-variation statistic:
@@
Following the recommendation of Cont and Das, (2024), we set $K_{\text{opt}}=\lfloor\sqrt{N}\rfloor$, where $N$ is the number of realised volatility observations for the given year and frequency. The corresponding block size is set to $n_{\text{opt}}=K_{\text{opt}}$, ensuring an equal partition. This results in $L=K_{\text{opt}}\times n_{\text{opt}}$ data points used in the estimation. When necessary, a few observations are discarded to make $L$ exactly divisible by $K_{\text{opt}}$. Table 2 reports the resulting partition parameters for each year and frequency.

### v3

# Multifractality in Bitcoin Realized Volatility: Implications for Rough Volatility Modelling

@@
Multifractality in Bitcoin Realized Volatility: Implications for Rough Volatility Modelling
@@
September 28, 2025
We assess the applicability of rough volatility models to Bitcoin realized volatility using the normalised $p$-variation framework of Cont and Das, (2024). Applying this model-free estimator to high-frequency Bitcoin data from 2017 to 2024 across multiple sampling resolutions, we find that the normalised statistic remains strictly negative, precluding the estimation of a valid roughness index. Stationarity tests and robustness checks reveal no significant evidence of non-stationarity or structural breaks as explanatory factors. Instead, convergent evidence from three complementary diagnostics, namely Multifractal Detrended Fluctuation Analysis, log-log moment scaling, and wavelet leaders, reveals a multifractal structure in Bitcoin volatility. This behaviour violates the homogeneity assumptions underlying rough volatility estimation and accounts for the estimator’s systematic failure. These findings suggest that while rough volatility models perform well in traditional markets, they are structurally misaligned with the empirical features of Bitcoin volatility.
@@ 1.1 Background and Motivation
The rough volatility framework has emerged as a prominent approach for modelling the fine-scale dynamics of financial market volatility. This development is grounded in empirical findings that indicate that realized volatility paths possess low Hölder regularity, which implies irregularity in their sample paths at high frequencies. Rough stochastic volatility models have been shown to replicate empirical features commonly observed in financial markets, including volatility clustering, short-term persistence, and the skew and term structure of implied volatility surfaces in equity and index markets Gatheral et al., (2018); Fukasawa et al., (2022). These observations have motivated attempts to extend the applicability of rough volatility models to nontraditional asset classes.
However, the extent to which this framework generalises to structurally distinct markets remains an open empirical question. Bitcoin, the most liquid and actively traded cryptocurrency, offers a particularly informative out-of-sample test. Unlike traditional equity or FX markets, Bitcoin exhibits fragmented liquidity across exchanges Makarov and Schoar, (2020), frequent discontinuities due to jumps Lahmiri and Bekiros, (2020), regime-switching volatility Caporale and Zekokh, (2021), and an evolving market microstructure shaped by protocol changes and shifting incentives Easley et al., (2019). This motivates an empirical examination of whether rough volatility assumptions hold in such a market.
@@
Prior applications of rough volatility methodologies to Bitcoin have predominantly employed parametric inference techniques or heuristic scaling diagnostics, often without systematically assessing assumptions of homogeneous Hölder continuity Bianchi et al., (2022). The normalised $p$-variation estimator recently introduced by Cont and Das, (2024) provides a model-free approach to estimating the pathwise roughness of volatility signals. This estimator offers two notable advantages over traditional methods based on log-volatility increments or variogram regressions: (i) it exhibits improved asymptotic properties under general conditions, and (ii) it is less sensitive to microstructure noise and finite-sample distortions. These properties make it especially suitable for high-frequency volatility data in contexts like cryptocurrency markets, where noise and irregular sampling are common.
This study builds on the normalised $p$-variation estimator introduced by Cont and Das, (2024) to perform a systematic evaluation of the roughness of Bitcoin realized volatility across a range of sampling frequencies and time horizons. This analysis examines the applicability of rough volatility models to Bitcoin and contributes to a deeper understanding of the structural features that differentiate cryptocurrency volatility dynamics.
@@ 1.2 Research Gap
Bitcoin’s volatility displays structural features such as multiscale clustering, episodic jumps, and irregular liquidity that differ markedly from the conditions under which rough volatility models have been validated.For instance, Bianchi et al., (2022) report low Hurst exponents for Bitcoin by fitting rough stochastic volatility models and performing scaling regressions on realized volatility. However, their approach implicitly assumes a global scaling regime and stationarity, without subjecting these conditions to empirical verification. These assumptions are critical, and their violation under multifractal dynamics or structural irregularities undermines the reliability of such roughness estimates. Crucially, no prior work has systematically examined whether Bitcoin realized volatility meets the homogeneity and stationarity conditions required for consistent estimation via the normalised $p$-variation method.
Moreover, recent research indicates that the observed roughness of realized volatility may arise from estimation error rather than genuine features of the spot volatility process Cont and Das, (2024); Rogers, (2019). In the case of Bitcoin, where multifractality and non-stationarity are theoretically plausible features, this raises serious doubts about the robustness and interpretive value of roughness estimates.
To our knowledge, no prior empirical study has systematically applied the $p$-variation diagnostic to Bitcoin realized volatility, nor has tested its structural compatibility with the assumptions underlying rough volatility models. It remains empirically unestablished whether Bitcoin volatility admits a well-defined roughness exponent. This study addresses this gap by applying the model-free normalised $p$-variation estimator alongside complementary diagnostic methods.
@@ 1.3 Research Question and Empirical Puzzle
This study tests whether Bitcoin realized volatility satisfies the homogeneous scaling assumptions necessary for roughness estimation under the rough volatility paradigm.
We investigate this question by applying the normalised $p$-variation methodology of Cont and Das, (2024) to Bitcoin realized volatility across multiple years and sampling frequencies. Preliminary results indicate that, in contrast to equity and foreign exchange markets, Bitcoin volatility does not appear to admit a well-defined roughness index. This stands in contrast to prior findings for equity and foreign exchange markets, where roughness estimates have exhibited consistent behaviour and robust interpretability.
@@ 1.5 Contributions
First, it offers an extensive application of the normalised $p$-variation framework to Bitcoin realized volatility across multiple years and sampling frequencies, making the analysis more comprehensive than any prior roughness estimation in cryptocurrency markets. The analysis reveals a systematic and robust failure of roughness estimation across all empirical configurations.
@@
Third, the results contribute to the evaluation of the rough volatility paradigm by showing that, while effective in equity and foreign exchange markets, rough stochastic volatility models face structural challenges when applied to Bitcoin realized volatility. This finding underscores the need for models that explicitly incorporate multifractal and scale-dependent dynamics to capture the unique characteristics of cryptocurrency markets. Examples of such alternatives include the Multifractal Random Walk Bacry et al., (2001) and Multifractal Stochastic Volatility Calvet and Fisher, (2002) models.
@@ 2.1 Rough Volatility and the Hurst Paradigm
Empirical support for the rough volatility hypothesis has emerged from multiple methodological approaches. These include the log-regression of realized volatility increments Gatheral et al., (2018), parametric estimation of fractional stochastic volatility models Fukasawa et al., (2022), and non-parametric diagnostics based on $p$-variation Cont and Das, (2024). Across these methods, estimated Hurst exponents in equity and index markets typically fall in the range 0.1 to 0.3. Rough models have demonstrated empirical success in capturing the fine-scale behaviour of implied volatility surfaces and short-term volatility dynamics.
The applicability of the rough volatility paradigm across asset classes remains contested. Rogers, (2019) and Cont and Das, (2024) have emphasised that low roughness estimates derived from realized volatility may not reflect intrinsic properties of the spot volatility process, but instead result from structural features of the data or estimation artefacts. For example, Rogers, (2019) showed that microstructure noise or irregular sampling can generate artificially low Hurst exponents, cautioning against interpreting realized path roughness as direct evidence of true volatility roughness. These confounding effects, including multifractal scaling, may induce apparent roughness even when the underlying volatility process is not genuinely rough. This underscores the need for model-free diagnostics when extending rough volatility models beyond traditional markets.
@@ 2.2 The Cont–Das Normalised $p$-Variation Estimator
To address the limitations of traditional approaches, we adopt the framework introduced by Cont and Das, (2024). This framework provides a model-free, pathwise-consistent estimator of roughness based on the scaling behaviour of the normalised $p$-variation statistic $W(L,K,p)$, computed over a sequence of block partitions of the observed time series. $W(L,K,p)$ measures how the $p$-th variation of the process accumulates across blocks, normalised in such a way that for a self-similar process with a global scaling law, the statistic equals one exactly at the correct scaling exponent $p^{*}$. Thus, plotting $\log W(L,K,p)$ as a function of $1/p$, one expects a zero-crossing at a unique value $p=p^{*}$, which defines the variation index and yields the corresponding roughness index $H=1/p^{*}$. The existence of this zero-crossing serves as a rigorous criterion for determining whether the process admits a global monofractal scaling regime consistent with Hölder-type roughness. For reference, the formal definition of $W(L,K,p)$ is provided in Section 3.1.
The normalised $p$-variation estimator is non-parametric, robust to non-Gaussian features, and directly applicable to realized volatility series. It serves not only as a roughness estimator but also as a structural diagnostic of the underlying process. Specifically, the presence or absence of a zero-crossing in $\log W(L,K,p)$ determines whether the data admit a well-defined global scaling law consistent with the rough volatility paradigm. If no zero-crossing is observed, it implies that no single roughness exponent can characterise the process, indicating a fundamental incompatibility with monofractal scaling assumptions.
Recent applications of the method to equity markets have validated its efficacy Cont and Das, (2024), producing roughness estimates consistent with the existing literature. In particular, Cont and Das apply the normalised $p$-variation estimator to 5-minute realized volatility of the S&P 500 using the Oxford-Man Institute dataset. They observe a well-defined zero-crossing in $\log W(L,K,p)$, with the resulting roughness index estimated in the range $H\in[0.05,0.25]$, consistent with prior findings from log-regression methods.
@@ 2.3 Multifractality in Financial Time Series
Empirical research has established that financial time series often exhibit multifractal properties. Foundational models such as the Multifractal Random Walk Bacry et al., (2001) and the Multifractal Stochastic Volatility model Calvet and Fisher, (2002) account for stylised features such as volatility clustering, heavy tails, and long memory through underlying multifractal dynamics. Subsequent studies have confirmed these properties across equities, foreign exchange, and commodities. More recent work has extended this evidence to cryptocurrencies. High-frequency Bitcoin returns and trading volume exhibit multifractal scaling when analysed using Multifractal Detrended Fluctuation Analysis Mensi et al., (2019); Matteo, (2007), and complex multiscale dependence has been observed in cross-correlation structures Watorek et al., (2022). Additional research shows that Bitcoin realized volatility also displays a multifractal structure, though weaker than that of returns Zhou et al., (2022). These findings support the view that multifractality is a robust and persistent feature of financial markets and offer a plausible explanation for the structural deviations from monofractal scaling observed in Bitcoin volatility.
@@ 3.2 Data Description and Preprocessing
A 90% completeness threshold was imposed at each frequency. For 1-minute and 5-minute data, all years from 2017 to 2024 met this criterion. For 10-minute data, 2022 and 2023 were excluded, while for 15-minute data, 2017, 2022, and 2023 were excluded. The remaining data were sufficient to observe clear patterns, and no contradictory behaviour was seen in partially excluded years at other frequencies. The complete list of retained and excluded years is given in Appendix Table A.1.
@@
Temporal structure and integrity of each retained window were further assessed. Large temporal gaps, defined as sequences of missing data exceeding one hour, were systematically identified at each frequency. Full gap statistics are provided in Appendix Table A.2.
@@
3.3 Realized Volatility Construction
@@
This absolute return series preserves the temporal resolution of the data and captures the local magnitude of return fluctuations. For notational convenience, we refer to $\{RV_{t}\}$ as the realized volatility series, although it is more accurately a proxy for instantaneous volatility rather than an integrated measure.
@@ 3.4 Diagnostic Exploration: Roughness vs $K$
Prior to finalising the roughness estimation procedure, we conduct a diagnostic analysis to examine how the estimated roughness $H_{L,K}$ depends on the number of blocks $K$ (which controls the partition size). The diagnostic analysis evaluates the behaviour of the roughness estimator $\widehat{H}^{L,K}$ as a function of the number of blocks $K$.
@@
To mitigate these issues, Cont and Das, (2024) recommend selecting $K$ such that $K\approx\sqrt{L}$, where $L$ is the number of available observations. This choice ensures a trade-off between bias and variance:
@@
Following the recommendation of Cont and Das, (2024), we set $K_{\text{opt}}=\lfloor\sqrt{N}\rfloor$, where $N$ is the number of realized volatility observations for the given year and frequency. The corresponding block size is set to $n_{\text{opt}}=K_{\text{opt}}$, ensuring an equal partition. This results in $L=K_{\text{opt}}\times n_{\text{opt}}$ data points used in the estimation. When necessary, a few observations are discarded to make $L$ exactly divisible by $K_{\text{opt}}$. Table 2 reports the resulting partition parameters for each year and frequency.